OpenStudy (solomonzelman):

This is very short, or at least supposed to be so. I saw a similar question and didn't know how to do it.

OpenStudy (solomonzelman):

\[4~Tan~(3x)\]in terms of tan x

OpenStudy (kc_kennylau):

Don't tell me you don't know how to type LaTeX, judging from your SS

OpenStudy (solomonzelman):

what do you mean? I did use latex for 4 tan 3x.

OpenStudy (kc_kennylau):

Did you use the "equation" button?

OpenStudy (solomonzelman):

Yes, don't you see that?

OpenStudy (kc_kennylau):

exactly, that's what i said

OpenStudy (solomonzelman):

\[\huge\color{red}{ 4~Tan~(3x) }~~~~i n~~~terms~~of\]\[\huge\color{blue}{ Tan~x }\]also I am 2 months here, so not that much.

OpenStudy (kc_kennylau):

ok let's go back to \(4\tan(3x)\). \[4\tan(3x)=4\tan(2x+x)\\\hspace{40pt}=4\frac{\tan2x+\tan x}{1-\tan2x\tan x}\\\hspace{40pt}=4\frac{\tan(x+x)+\tan x}{1-\tan(x+x)\tan x}\\\hspace{40pt}=4\frac{\dfrac{2\tan x}{1-\tan^2x}+\tan x}{1-\dfrac{2\tan x}{1-\tan^2x}\tan x}\]

OpenStudy (kc_kennylau):

Lemme introduce to you eight functions in \(\LaTeX\):)

OpenStudy (kc_kennylau):

\LaTeX \cos \sin \tan \sec \csc \cot \mbox{abcd}

OpenStudy (solomonzelman):

\(\huge\color{blue}{\huge {\bbox[5pt,cyan,border:2px solid purple]{Alright.}}}\)

OpenStudy (solomonzelman):

factorize them what do you mean?

OpenStudy (kc_kennylau):

I believe you can simplify it then? :)

OpenStudy (solomonzelman):

what the thing above?

OpenStudy (kc_kennylau):

exactly

OpenStudy (solomonzelman):

Sometimes, I cope paste and post replies, dk how, sorry...

OpenStudy (solomonzelman):